OpenStudy (chillhill):

The equation kx^2 + 3x + 1=0 has exactly one solution. Find the value of k.

OpenStudy (michele_laino):

we have to require that the discriminat of such trinomial, is zero, namely: \[\huge \Delta = 9 - 4k = 0\]

jhonyy9 (jhonyy9):

so you need to know when in what case of discriminant one quadratic has one solution ?

OpenStudy (whpalmer4):

standard form of a quadratic and its solution are \[ax^2 + bx + c = 0\]\[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]to have only one solution, the term under the radical (aka the discriminant) must be \(0\) Viewed a different way, the vertex of the parabola will be on the x-axis, which is the only way a quadratic can have a single solution. with our equation, we have \(a=k\) and \(b=3\), so the value of \(x\) that solves the equation will be \[x = -\frac{3}{2k}\] Note that this is NOT the answer to your question! However, it can be used to check your answer for correctness.